An iron bar of mass, 0.57 kg, slides down a frictionless incline 0.66 m high.  At the bottom of the incline it collides and sticks to a magnet of mass, 1.28 kg.  What is the speed of the combined masses after collision.
Answer #1

here ,

h = 0.66 m

let the speed of iron bar at the bottom is u m/s

Now , Using conservation of energy

0.5 * m*u^2 = m*g*h

u = sqrt(2*0.66*9.8)

u = 3.6 m/s

Now , let speed of combined masses is v m/s

Using conservation of momentum

3.6 * 0.57 = (0.57 + 1.28)* v

v = 1.11 m/s

the speed of combined masses is 1.11 m/s
Answer #2

As the bar moves 0.66 meter down, its kinetic energy increases as its potential energy decreases.
KE =